OpenStudy (ashking1):

What are the solutions of x2 + 5x - 24 = 0? Show all work for full credit.

OpenStudy (lgbasallote):

you can factor that out...but if you can't just use quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] \[x = \frac{-(5) \pm \sqrt{5^2 - 4(1)(-24)}}{2(1)}\]

OpenStudy (anonymous):

x-3 x+8

OpenStudy (lgbasallote):

now if you simplify that... \[x = \frac{-5 \pm \sqrt{25 + 96}}{2}\] \[x = \frac{-5 \pm \sqrt{121}}{2}\] \[x = \frac{-5 \pm 11}{2}\] \[x = \frac{-5 - 11}{2} ; x = \frac{-5 + 11}{2}\]

OpenStudy (anonymous):

x^2 + 5x - 24 = 0 (x+8)(x-3) =0 x = -8 or x = 5

OpenStudy (lgbasallote):

simplifying further... \[x = \frac{-16}{2}; x = \frac{6}{2}\]

OpenStudy (lgbasallote):

got it @ashking1